# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
(A3.5) Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), and (A3) hold for all ETTs with at most $n$ rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ could be extended further with $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n})\setminus\{\gamma_{n+1}\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then the coloring $\varphi^{\prime}=\varphi/P$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
@@
Let $T_{n+1}$, $\alpha$, $\beta$, and $P$ be as described above. So $\varphi^{\prime}=\varphi/P$ is $(T_{n+1},D_{n+1},\varphi)$-stable. So Lemma 4.7 follows immediately from (A3) and (A2.5) if $T_{n+1}$ is $RE$ finished. Thus, we assume that $\Theta_{n+1}=RE$ with initial index $h$ such that $1\leq h\leq n$. In this case, following (A2.5), we only need to confirm that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, where $\gamma_{h}=\gamma_{n+1}$ and $\delta_{h}=\delta_{n+1}$. Note that $\delta_{h}$ is a defective color of $T_{n+1}$ under $\varphi$, because $\Theta_{n+1}=RE$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. So $\alpha\neq\delta_{h}$ and $\alpha\neq\gamma_{h}$. Since $T_{h}$ is elementary by (A1) and $\overline{\varphi}(T_{1})\geq 5$ by Lemma 3.2, there exists a color $\eta\in\overline{\varphi}(T_{h})\setminus\{\gamma_{h},\alpha\}$. Moreover, $\eta\neq\delta_{h}$ as $\delta_{h}$ is a defective color of $T_{n+1}$ under $\varphi$. Thus $\{\alpha,\eta\}\cap\{\gamma_{h},\delta_{h}\}=\emptyset$. Let $\pi=(G-T_{n+1},\alpha,\eta)$. Clearly, $\pi$ is $(T_{n+1},D_{n+1},\varphi)$-stable and $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$. So $\pi$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring. Now $P$ is a $(\eta,\beta)$-path not intersecting $T_{h}$. So by (A3) on $T_{h}$ and $P$, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi/P$ as under $\pi$ and $\varphi$. However, since $\{\alpha,\eta\}\cap\{\gamma_{h},\delta_{h}\}=\emptyset$ and $P$ does not intersect $T_{n+1}$, the $\gamma_{h}$ colored edges and $\delta_{h}$ colored edges are the same under $\varphi^{\prime}=\varphi/P$ as under $\pi/P$. So $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\pi/P$ and $\varphi$, as desired. ∎
@@ 5 Elementary
In this section, we will prove the elementariness of $T$ if $T$ is an ETT with ladder $T_{0}\subset T_{1}\subset\dots\subset T_{n}\subset T$ under a $k$-triple $(G,e,\varphi)$ (A1). In the proof we will divide the sequence of $T-T_{n}$ into a number of subsequence. We call the nested sequence
@@
A hierarchy (5.1) of $T$ is called good under $\varphi$ if for all $j$ with $0\leq j\leq q$ and all $\delta_{m}\in D_{n,j}$, there exists a $2$-color subset $\Gamma^{j}_{m}=\{\gamma^{j}_{m_{1}},\gamma^{j}_{m_{2}}\}\subseteq\overline{\varphi}(T_{n,j})\setminus\{\gamma_{n}\,\gamma,\tau\}$ with $\{\gamma,\tau\}\subseteq\overline{\varphi}(x_{e})$ and $\gamma\neq\tau$, such that
@@
The case $\gamma_{n}\in\{\gamma,\tau\}$ could happen. The sets $\Gamma^{j}_{m}$ are referred to as $\Gamma$-sets for the hierarchy of $T$ under $\varphi$, and let $\Gamma^{j}=\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j}_{m}$. We say a hierarchy of $T=T_{n,q+1}$ is good up to itself under $\varphi$ if $T_{n,q+1}$ is $(\cup_{\delta_{m}\in D_{n,q+1}}\Gamma^{q}_{m})^{-}$-closed. For convenience, we say $T_{n}$ has a good hierarchy of $(-1)$-levels up to itself under $\varphi$.
@@
(A1) Let $n$ be a positive integer. Suppose (A2), (A2.5),(A3), and (A3.5) hold for ETTs with at most $n$ rungs, and (A1) holds for ETTs satisfying MP with at most $(n-1)$-rungs. If $T$ is an ETT with ladder $T_{1}\subset T_{2}\subset\dots\subset T_{n}\subset T_{n}\subset T\subseteq T_{n+1}$ under $\varphi$, then the following statements A and B hold, which imply $T$ is elementary.
@@
    If A holds, then under $\varphi$, there exists a closed ETT $T^{\prime}$ having a good hierarchy with $V(T_{n+1})=V(T^{\prime})$ and ladder $T_{1}\subset T_{2}\subset\dots\subset T_{n}\subset T_{n}\subset T^{\prime}$.
@@
We prove statement A by induction on $q$ which is the number of levels. Denote $T$ by $T_{1}\subset T_{2}\subset\dots\subset T_{n}:=T_{n,0}\subset T_{n,1}\subset\dots\subset T_{n,q}\subset T:=T_{n,q+1}$. When $q=0$, we have $T_{n,q}=T_{n,0}=T_{n}$. Since $T_{n}$ is an ETT under $\varphi_{n-1}$ with $n(T_{n})=n-1$, it is elementary by (A1) with $(n-1)$-rungs. Now we assume $T_{n,q}$ is elementary and show $T_{n,q+1}=T$ is. Denote $T$ by $\{T_{n,q},e_{1},y_{1},e_{2},...,e_{p},y_{p}\}$ following the order $\prec_{L}$. We define the path number $p(T)$ of $T$ as the smallest index $i\in{1,...,p}$ such that the sequence $y_{i}T:=(y_{i},e_{i+1},...,e_{p},y_{p})$ corresponds to path in $G$. Suppose on the contrary that $T$ is a counterexample to the theorem, i.e., $T$ has a good hierarchy with $q$-levels under $\varphi$, but $V(T)$ is not elementary. Furthermore, we assume that among all counterexamples, the following two conditions hold:
@@ 5.2.1 A few basic properties
Let $j$ be the smallest index such that $\alpha\in\overline{\varphi}(T_{n,j})$. If $j=0$, then $\alpha$ is closed for $T_{n,0}=T_{n}$. In the case that $j\geq 1$, we have $\alpha\notin\overline{\varphi}(T_{n,j-1})$. By Definition 4, $\alpha\notin\Gamma^{j-1}$, and therefore $\alpha$ is closed for $T_{n,j}$ by Definition 4 (iv). So there exists $r$ as the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$. The rest of the current claim holds trivially if $r=q$, so we may assume $r<q$. Since $r$ is the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$, for each $k$ with $r<k\leq q$, we have $\alpha\in\cup_{\delta_{m}\in D_{n,k}}\Gamma^{k-1}_{m}$ by Definition 4 (iv). So following Definition 4 (i), we see that $\alpha\notin\varphi(T_{n,k}-T_{n,k-1})$, and therefore $\alpha\notin\varphi(T_{n,q}-T_{n,r})$. ∎
@@
Since $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, any closure of $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\varphi_{\pi}$. Because $\pi$ is also $(T_{n,j},D_{n},\varphi)$-stable, $T_{n,j}$ remains a closure of $T_{n}+f_{n}$ under $\varphi_{\pi}$, and therefore is an ETT with $n$-rungs. The hierarchy property and the fact that $T_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed when $j\geq 1$ follow from Definition 4 of $T_{n,q}$. ∎
@@
Let $T$ be an ETT under $\varphi$ with $n$-rungs and a good hierarchy of $j$-levels with $j\leq q$, and let $\pi$ be $T_{n}+f_{n}$ mod $\varphi$. It is worth pointing out that under $\pi$, if $T_{n,j}$ has a good hierarchy up to itself of $(j-1)$-levels and $T$ is still an ETT, then to check whether $T$ or any tree-sequence obtained from $T$ by TAA has a good hierarchy of $j$ levels under $\pi$, we only need to check Definition 4 (i) for $T_{n,j+1}$ which is $T$ or a tree-sequence obtained from $T_{n,j}$ by TAA, because Definition 4 (ii), (iii) and (iv) do not involve $T_{n,j+1}$. So if $\pi$ is also $(T,D_{n},\varphi)$-weakly stable, then following Lemma 3.1, $T$ also has a good hierarchy, as it satisfies Definition 4 (i) under $\pi$ the same way as under $\varphi$.
@@
We first prove (a). Since $\gamma_{n},\gamma,\tau\in\overline{\varphi}(T_{n})$ and $T_{n}$ is closed, the colors $\gamma$, $\tau$ and $\gamma_{n}$ are all closed for $T_{n,0}$ under $\varphi$. Moreover, $\gamma_{n},\gamma,\tau\notin\Gamma_{j}$ by the choice of $\Gamma$ sets in Definition 4. Then for each $1\leq i\leq j$, since $T_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed following Definition 4 (iv) and (1), all three colors $\gamma$, $\tau$ and $\gamma_{n}$ are closed for $T_{i}$ under $\varphi$ as well. We then prove (b). The proof of (b) is actually very similar to the proof of interchangeability from Lemma 4.1 (A2). Assume on the contrary: Let $j$ be the smallest index such that there exist two $(\alpha,\beta)$-paths $Q_{1}$ and $Q_{2}$ intersecting the ETT $T_{n,j}$ as described above, where $\alpha\in\overline{\varphi}(T_{n,j})\setminus\{\gamma_{n}\}$ and is closed in $T_{n,j}$. Since $T_{n,j}$ is elementary by induction hypothesis, paths $Q_{1}$ and $Q_{2}$ together yield at least two $(\alpha,\beta)$-exit paths, say $P_{1}$ with exit vertex $b_{1}$ and $P_{2}$ with exit vertex $b_{2}$, respectively. Renaming subscripts if necessary, we may assume that $b_{1}\prec b_{2}$ along $T_{n,j}$. Note that we have $j\geq 1$, because the case that $j=0$ follows directly from (A2) on $T_{n}=T_{n,0}$. Let $h$ be the initial index for $\Theta_{n+1}$. So $\gamma_{h}=\gamma_{n}$. For any color $\pi\in\overline{\varphi}(T_{n,j})$ that is closed for $T_{n,j}$ with $\pi\neq\gamma_{h}$, let $\varphi_{\pi}=\varphi/(G-T_{n,j},\alpha,\pi)$. We claim that (2) $\varphi_{\pi}$ is $T_{n}+f_{n}$ mod $\varphi$, and $T_{n,j}$ is an ETT with $n$-rungs under $\varphi_{\pi}$ that has a good hierarchy up to itself. Since $\alpha$ and $\pi$ are closed for $T_{n,j}$, $\varphi_{\pi}$ is $(T_{n,j},D_{n},\varphi)$-stable. Note that $\pi$ and $\alpha$ are different from $\gamma_{h}=\gamma_{n}$. So if they are all different from $\delta_{h}$, clearly $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\pi}$ as $\varphi$. If one of them is $\delta_{h}$, then since $\alpha$, $\pi$ and $\gamma_{h}$ (see (a)) are closed for $T_{n,j}$, all the $(\gamma_{h},\delta_{h})$-chains intersecting $T_{h}$ are contained in $G[V(T_{n,j})]$, so $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\pi}$ as $\varphi$. Thus either way, $\varphi_{\pi}$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) because $\varphi_{\pi}$ is also $(T_{n},D_{n},\varphi)$-stable. We then have (1) from Claim 5.2.
@@
Without loss of generality, we assume one of $\gamma$ and $\tau$, say $\gamma$, is different from $\gamma_{h}$. We claim that (3) $b_{2}\notin V(T_{n,j-1})$, and $P_{2}$ is a subpath of a $(\gamma,\beta)$-path $P$ not intersecting $T_{n,j-1}$ under $\varphi_{\gamma}$. By (a) and (2), $\varphi_{\gamma}=\varphi/(G-T_{n,j},\alpha,\gamma)$ is $T_{n}+f_{n}$ mod $\varphi$, and $T_{n,j}$ remains an ETT with a good hierarchy under $\varphi_{\gamma}$. Since $\gamma\neq\gamma_{h}$, we see that there is at most one $(\gamma,\beta)$-path intersecting $T_{n,j-1}$ by the minimality of $j$. Because $P_{1}$ and $P_{2}$ are $(\gamma,\beta)$-exit paths of $T_{n,j}$ under $\varphi_{\gamma}$ with exit vertices $b_{1}\prec b_{2}$, we see that $P_{2}$ is a subpath of a $(\gamma,\beta)$-path $P$ not intersecting $T_{n,j-1}$ under $\varphi_{\gamma}$, and thus $b_{2}\notin T_{n,j-1}$.
@@
Case I:
$\beta\neq\delta_{h}$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
(A3.5) Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), and (A3) hold for all ETTs with at most $n$ rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ could be extended further with $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma_{n+1}\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then the coloring $\varphi^{\prime}=\varphi/P$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
@@
Let $T_{n+1}$, $\alpha$, $\beta$, and $P$ be as described above. So $\varphi^{\prime}=\varphi/P$ is $(T_{n+1},D_{n+1},\varphi)$-stable. So Lemma 4.7 follows immediately from (A3) and (A2.5) if $T_{n+1}$ is $\operatorname{RE}$ finished. Thus, we assume that $\Theta_{n+1}=\operatorname{RE}$ with initial index $h$ such that $1\leq h\leq n$. In this case, following (A2.5), we only need to confirm that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, where $\gamma_{h}=\gamma_{n+1}$ and $\delta_{h}=\delta_{n+1}$.
Note that $\delta_{h}$ is a defective color of $T_{n+1}$ under $\varphi$, because $\Theta_{n+1}=\operatorname{RE}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$. So $\alpha\neq\delta_{h}$ and $\alpha\neq\gamma_{h}$. Since $T_{h}$ is elementary by (A1) and $\overline{\varphi}(T_{1})\geq 5$ by Lemma 3.2, there exists a color $\eta\in\overline{\varphi}(T_{h})\setminus\{\gamma_{h},\alpha\}$. Moreover, $\eta\neq\delta_{h}$ as $\delta_{h}$ is a defective color of $T_{n+1}$ under $\varphi$. Thus $\{\alpha,\eta\}\cap\{\gamma_{h},\delta_{h}\}=\emptyset$. Let $\pi=\varphi/(G-T_{n+1},\alpha,\eta)$. Clearly, $\pi$ is $(T_{n+1},D_{n+1},\varphi)$-stable and $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$. So $\pi$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring. Now $P$ is an $(\eta,\beta)$-path not intersecting $T_{h}$. So by (A3) on $T_{h}$ and $P$, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi/P$ as under $\pi$ and $\varphi$. However, since $\{\alpha,\eta\}\cap\{\gamma_{h},\delta_{h}\}=\emptyset$ and $P$ does not intersect $T_{n+1}$, the $\gamma_{h}$ colored edges and $\delta_{h}$ colored edges are the same under $\varphi^{\prime}=\varphi/P$ as under $\pi/P$. So $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\pi/P$ and $\varphi$, as desired.
@@ 5 Elementary
In this section, we will prove the elementariness of $T$ if $T$ is an ETT with ladder $T_{0}\subset T_{1}\subset\dots\subset T_{n}\subset T$ under a $k$-triple $(G,e,\varphi)$ (A1). Let $T-T_{n}$ be the subsequence of $T$ after removing $T_{n}$. In the proof we will divide $T-T_{n}$ into a number of subsequences. We call the nested sequence
@@
A hierarchy (5.1) of $T$ is called good under $\varphi$ if for all $j$ with $0\leq j\leq q$ and all $\delta_{m}\in D_{n,j}$, there exists a $2$-color subset $\Gamma^{j}_{m}=\{\gamma^{j}_{m_{1}},\gamma^{j}_{m_{2}}\}\subseteq\overline{\varphi}(T_{n,j})\setminus\{\gamma_{n},\gamma,\tau\}$ with $\{\gamma,\tau\}\subseteq\overline{\varphi}(x_{e})$ and $\gamma\neq\tau$, such that
@@
The case $\gamma_{n}\in\{\gamma,\tau\}$ could happen. The sets $\Gamma^{j}_{m}$ are referred to as $\Gamma$-sets for the hierarchy of $T$ under $\varphi$, and let $\Gamma^{j}=\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j}_{m}$. We say a hierarchy of $T=T_{n,q+1}$ is good up to itself under $\varphi$ if $T_{n,q+1}$ is $(\cup_{\delta_{m}\in D_{n,q+1}}\Gamma^{q}_{m})^{-}$-closed. For convenience, we say $T_{n}$ has a good hierarchy of (1)(-1) levels up to itself under $\varphi$.
@@
(A1) Let $n$ be a positive integer. Suppose (A1), (A2), (A2.5),(A3), and (A3.5) hold for ETTs with at most n1n-1 rungs. If $T$ is an ETT with ladder $T_{1}\subset T_{2}\subset\dots\subset T_{n}\subset T\subseteq T_{n+1}$ under $\varphi$, then the following statements A and B hold, which imply $T$ is elementary.
@@
    If A holds, then under $\varphi$, there exists a closed ETT $T^{\prime}$ having a good hierarchy with $V(T_{n+1})=V(T^{\prime})$ and ladder $T_{1}\subset T_{2}\subset\dots\subset T_{n}\subset T^{\prime}$.
@@
We prove statement A by induction on $q$ which is the number of levels. Denote $T$ by $T_{1}\subset T_{2}\subset\dots\subset T_{n}:=T_{n,0}\subset T_{n,1}\subset\dots\subset T_{n,q}\subset T:=T_{n,q+1}$. When $q=0$, we have $T_{n,q}=T_{n,0}=T_{n}$. Since $T_{n}$ is an ETT under $\varphi_{n-1}$ with $n(T_{n})=n-1$, it is elementary by (A1) with (n1)(n-1)-rungs. Now we assume $T_{n,q}$ is elementary and show $T_{n,q+1}=T$ is. Denote $T$ by $\{T_{n,q},e_{1},y_{1},e_{2},...,e_{p},y_{p}\}$ following the order $\prec$. We define the path number $p(T)$ of $T$ as the smallest index $i\in{1,...,p}$ such that the sequence $y_{i}T:=(y_{i},e_{i+1},...,e_{p},y_{p})$ corresponds to path in $G$. Suppose on the contrary that $T$ is a counterexample to the theorem, i.e., $T$ has a good hierarchy with $q$ levels under $\varphi$, but $V(T)$ is not elementary. Furthermore, we assume that among all counterexamples, the following two conditions hold:
@@ 5.2.1 A few basic properties
Let $j$ be the smallest index such that $\alpha\in\overline{\varphi}(T_{n,j})$. If $j=0$, then $\alpha$ is closed for $T_{n,0}=T_{n}$. In the case that $j\geq 1$, we have $\alpha\notin\overline{\varphi}(T_{n,j-1})$. By Definition 4, $\alpha\notin\Gamma^{j-1}$, and therefore $\alpha$ is closed for $T_{n,j}$ by Definition 4 (iv). So there exists $r$ as the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$. The rest of the current claim holds trivially if $r=q$, so we may assume $r<q$. Since $r$ is the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$, for each $k$ with $r<k\leq q$, we have $\alpha\in\cup_{\delta_{m}\in D_{n,k}}\Gamma^{k-1}_{m}$ by Definition 4 (iv). So following Definition 4 (i), we see that $\alpha\notin\varphi(T_{n,k}-T_{n,k-1})$, and therefore $\alpha\notin\varphi(T_{n,q}-T_{n,r})$.
@@
Since $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, any closure of $T_{n}+f_{n}$ is an ETT with $n$-rungs under $\pi$. Because $\pi$ is also $(T_{n,j},D_{n},\varphi)$-stable, $T_{n,j}$ remains a closure of $T_{n}+f_{n}$ under $\pi$, and therefore is an ETT with $n$-rungs. The hierarchy property and the fact that $T_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed when $j\geq 1$ follow from Definition 4 of $T_{n,q}$. ∎
@@
Let $T$ be an ETT under $\varphi$ with $n$-rungs and a good hierarchy of $j$ levels with $j\leq q$, and let $\pi$ be $T_{n}+f_{n}$ mod $\varphi$. It is worth pointing out that under $\pi$, if $T_{n,j}$ has a good hierarchy up to itself of (j1)(j-1) levels and $T$ is still an ETT, then to check whether $T$ or any tree-sequence obtained from $T$ by TAA has a good hierarchy of $j$ levels under $\pi$, we only need to check Definition 4 (i) for $T_{n,j+1}$ which is $T$ or a tree-sequence obtained from $T_{n,j}$ by TAA, because Definition 4 (ii), (iii) and (iv) do not involve $T_{n,j+1}$. So if $\pi$ is also $(T,D_{n},\varphi)$-weakly stable, then following Lemma 3.1, $T$ also has a good hierarchy, as it satisfies Definition 4 (i) under $\pi$ the same way as under $\varphi$.
@@
We first prove (a). Since $\gamma_{n},\gamma,\tau\in\overline{\varphi}(T_{n})$ and $T_{n}$ is closed, the colors $\gamma$, $\tau$ and $\gamma_{n}$ are all closed for $T_{n,0}$ under $\varphi$. Moreover, $\gamma_{n},\gamma,\tau\notin\Gamma_{j}$ by the choice of $\Gamma$ sets in Definition 4. Then for each $1\leq i\leq j$, since $T_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed following Definition 4 (iv) and (1), all three colors $\gamma$, $\tau$ and $\gamma_{n}$ are closed for $T_{i}$ under $\varphi$ as well.
@@
We then prove (b). The proof of (b) is actually very similar to the proof of interchangeability from Lemma 4.1 (A2). Assume on the contrary: Let $j$ be the smallest index such that there exist two $(\alpha,\beta)$-paths $Q_{1}$ and $Q_{2}$ intersecting the ETT $T_{n,j}$ as described above, where $\alpha\in\overline{\varphi}(T_{n,j})\setminus\{\gamma_{n}\}$ and is closed in $T_{n,j}$. Since $T_{n,j}$ is elementary by induction hypothesis, paths $Q_{1}$ and $Q_{2}$ together yield at least two $(\alpha,\beta)$-exit paths, say $P_{1}$ with exit vertex $b_{1}$ and $P_{2}$ with exit vertex $b_{2}$, respectively. Moreover, if $\beta\notin\overline{\varphi}(T_{n,j})$, we have an additional $(\alpha,\beta)$-exit path $P_{3}$ with exit vertex $b_{3}$. Renaming subscripts if necessary, we may assume that $b_{1}\prec b_{2}$ (and $b_{2}\prec b_{3}$ if $P_{3}$ exists) along $T_{n,j}$. Note that we have $j\geq 1$, because the case that $j=0$ follows directly from (A2) on $T_{n}=T_{n,0}$. Let $h$ be the initial index for $\Theta_{n+1}$. So $\gamma_{h}=\gamma_{n}$. For any color $\pi\in\overline{\varphi}(T_{n,j})$ that is closed for $T_{n,j}$ with $\pi\neq\gamma_{h}$, let $\varphi_{\pi}=\varphi/(G-T_{n,j},\alpha,\pi)$. We claim that
(2) $\varphi_{\pi}$ is $T_{n}+f_{n}$ mod $\varphi$, and $T_{n,j}$ is an ETT with $n$-rungs under $\varphi_{\pi}$ that has a good hierarchy up to itself.
Since $\alpha$ and $\pi$ are closed for $T_{n,j}$, $\varphi_{\pi}$ is $(T_{n,j},D_{n},\varphi)$-stable. Note that $\pi$ and $\alpha$ are different from $\gamma_{h}=\gamma_{n}$. So if they are all different from $\delta_{h}$, clearly $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\pi}$ as $\varphi$. If one of them is $\delta_{h}$, then since $\alpha$, $\pi$ and $\gamma_{h}$ (see (a)) are closed for $T_{n,j}$, all the $(\gamma_{h},\delta_{h})$-chains intersecting $T_{h}$ are contained in $G[V(T_{n,j})]$, so $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{\pi}$ as $\varphi$. Thus either way, $\varphi_{\pi}$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) because $\varphi_{\pi}$ is also $(T_{n},D_{n},\varphi)$-stable. We then have (1) from Claim 5.2.
Without loss of generality, we assume one of $\gamma$ and $\tau$, say $\gamma$, is different from $\gamma_{h}$. We claim that
(3) We may assume that $b_{2}\notin V(T_{n,j-1})$, and $P_{2}$ is a subpath of a $(\gamma,\beta)$-path $P$ not intersecting $T_{n,j-1}$ under $\varphi_{\gamma}$.
@@
By (a) and (2), $\varphi_{\gamma}=\varphi/(G-T_{n,j},\alpha,\gamma)$ is $T_{n}+f_{n}$ mod $\varphi$, and $T_{n,j}$ remains an ETT with a good hierarchy under $\varphi_{\gamma}$. Since $\gamma\neq\gamma_{h}$, we see that there is at most one $(\gamma,\beta)$-path intersecting $T_{n,j-1}$ by the minimality of $j$. If $\beta\in\overline{\varphi}(T_{n,j})$, then $P_{2}$ is a subpath of a $(\gamma,\beta)$-path $P$ not intersecting $T_{n,j-1}$ under $\varphi_{\gamma}$, and thus $b_{2}\notin T_{n,j-1}$ as desired. Now if $\beta\notin\overline{\varphi}(T_{n,j})$, then at least one of $P_{2}$ and $P_{3}$ ($P_{3}$ exists when $\beta\notin\overline{\varphi}(T_{n,j})$) is a subpath of a $(\gamma,\beta)$-path $P$ not intersecting $T_{n,j-1}$ under $\varphi_{\gamma}$. Relabel $P_{3}$ to $P_{2}$ and $b_{3}$ to $b_{2}$ if necessary, we have as claimed. Note that we will not use $b_{3}$ after this point, so $b_{3}\prec b_{2}$ after relabeling will not be an issue.
@@
Case I: $\beta\neq\delta_{h}$.